OpenStudy (anonymous):

Find an equation of the circle with center (18,-1) and radius 5 in the form of (x-A)^2+(y-B)^2=C where A, B, C are constant.

OpenStudy (anonymous):

\[(x-A)^2+(y-B)^2=C^2\] given that \(A=18\) \(B=-1\) \(C=5\) Plug those numbers in the equation.